OpenStudy (anonymous):

Im still confused on how I find the solution to the system of equations with this problem... what about use a matrix to find the solution to te system of equations? -8x-9y=-16 6x-9y=-108 (-6,8) (6,8) (8,-6) (6,-8

OpenStudy (anonymous):

You do not need to use matrices

OpenStudy (anonymous):

i dont know the answer, or how to do it?

OpenStudy (anonymous):

-8x-9y=-16 6x-9y=-108 Subtract the second from the first to get -14x =-16+108 -14 x =92 x=-46/7

OpenStudy (anonymous):

Replace in one of equations and get y=160/21

OpenStudy (anonymous):

All your choices are not correct or the equations are not typed correctly

OpenStudy (anonymous):

use matrix to find the solution to the system of equations? -8x-8y=-16 6x-9y=-108

OpenStudy (anonymous):

sorry i typed it wrong earlier

OpenStudy (anonymous):

theres a matrix called A and is equal to (-8 -8) (6 -9) then you have B which is (-16) (-108)

OpenStudy (anonymous):

first you find the determinant of the matrix A which is (-8)*(-9) - (-8)*6 and we'll call it detA

OpenStudy (anonymous):

now to find x you need to replace the first column of A with the column-matrix B and also calculate its determinant (the product of the main diagonal minus the product of the secondary diagonal)

OpenStudy (anonymous):

after youve done this, you simply do (the last determinant)/detA and you do the same for Y, only replacing the second column of A with the column-matrix B

OpenStudy (anonymous):

thats all. It's called the Cramer rule

OpenStudy (anonymous):

so whats the answer? (6,8)

OpenStudy (anonymous):

thats left for you to do :)

OpenStudy (anonymous):

not that hard really. simple multiplications